# Heuristic Adaptation of Potentially Misspecified Domain Support for Likelihood-Free Inference in Stochastic Dynamical Systems

### v1

@@ II-B Iterative posterior refinement
In addition to iterative posterior refinement, sequential inference can involve progressive densification of the dataset $\mathcal{D}$ used to train the CDF $q_{\phi}$ (alg. 1, 10[35]. In this way, our belief about $\boldsymbol{\theta}$ evolves recursively as more informative simulated evidence accumulates. When the samples drawn from each iteration’s posterior are aggregated in the same dataset, densifying certain areas of the supported parameter space, this proxies the reweighing of the prior to adaptively shift focus toward informative regions. Thus, in practice, we can ignore the proposal prior $\tilde{p}(\boldsymbol{\theta})$ and infer the posterior using only the (desirable) prior $p(\boldsymbol{\theta})$ and the CDF.
@@ III A demo of the support misspecification issue
We experiment with the stochastic dynamical benchmarks of the Lotka-Volterra and M/G/1 queue, as implemented for the seminal LFI work of [11]. The hyperparameters of the MDNN used to approximate the CDF (alg. 1, 12) in the following experiments are given in App. B, Table II.
@@ IV Heuristic support optimisation methods
We treat support expansion as window stretching [40] rather than window sliding [9]: instead of shifting the sampling window, we iteratively accumulate new parameter samples guided by the latest posterior (alg. 1, 10). We compactly denote a distribution $p$ with support $\Theta\subseteq\mathbb{R}^{(D)}$ as $p_{(\Theta)}$, but do this sparingly. This arbitrarily denotes that the distribution is supported by a $D$-dim. rectangular range defined by the bounds $\Theta=[\boldsymbol{\theta}_{\min}^{(d)},\boldsymbol{\theta}_{\max}^{(d)}]_{d=1}^{D}$.
@@ IV-A BayesSim-EDGE: expanding bounds via edge mass
EDGE is configurable through its hyperparameters. These are the edge mass threshold $\tau$, the edge zone fraction $\delta$, the edge expansion factor $\eta$, and the physically feasible domain limits $\Phi$. Specifically, $\delta$ is the size of the area within which we consider any accumulated probability mass to be near the edge. This is the sensitivity of our expansion criterion, which compares the accumulated probability mass with a threshold $\tau$. We denote the left and right edge zones ($\min$ and $\max$ side) as $\Delta_{L}^{(d)}$ and $\Delta_{R}^{(d)}$ for a dimension $d$ of $\boldsymbol{\theta}$, and the respective accumulated masses as $M_{L}^{(d)}$ and $M_{R}^{(d)}$. If $\tau$ has been exceeded and if an expansion would still be within $\Phi$ (20, LABEL: and 23), we consider the expansion criterion to be satisfied and expand the respective domain boundary by a factor of $\eta$ (21, LABEL: and 24). If $\tau$ has been exceeded, but the expansion would have been beyond the extremum of the respective $\Phi$ limit, then this becomes the new boundary.
@@
EDGE explores beyond the initial parameter ranges if the inferred posterior (12) confidently points toward the edges of the domain. This confidence is tuned through $\tau$ and $\delta$. Thus, within $N_{\text{LFI}}$ iterations, we can converge to an adapted support $\Theta_{N_{\text{LFI}}}$ in conjunction with a posterior $\hat{p}$.
@@ IV-B BayesSim-MODE: mode-shift-based expansion
Similarly to EDGE, the standard BayesSim assumptions apply. We also have MODE’s hyperparameters, which are the mode shift threshold $\nu_{\text{TH}}$, the bounds proximity threshold $\rho$, the MoG weights summary threshold $\tau$, the bounds expansion factor $\eta$, and the physically feasible domain limits $\Phi$. The main difference between MODE and EDGE is that we now calculate the shift of the MoG posterior modes $\boldsymbol{\mu}_{i}$ compared to the previous iteration $\boldsymbol{\mu}_{i-1}$ (21) along with their normalised proximity to the bounds (23). Thus, we do not check for $\Theta$ adaptation at the end of the first inference iteration. In all subsequent iterations, the MoG details of the previous iteration are accessible through the reference prior.
We accumulate the mixture coefficients (weights) of each mode that satisfies the shift and proximity thresholds for the left or right bound (26, LABEL: and 28) of a parameter $d$. The accumulated weights $W_{L}^{(d)}$, $W_{R}^{(d)}$ are used for the expansion criteria (31, LABEL: and 34). If $\tau$ has been exceeded and if an expansion would be within $\Phi$ (31, LABEL: and 34), we expand it by a factor of $\eta$ (32, LABEL: and 35). If $\tau$ has been exceeded, but the expansion would have been beyond the corresponding $\Phi$ limit, then this becomes the new boundary.
@@ IV-C BayesSim-CENTRE: centring-based adaptive bounds
Thus, beyond the standard BayesSim assumptions, we now only have physically feasible domain limits $\Phi$. In each inference iteration, following the posterior $\hat{p}$ update (10), we compute the weighed mean $\mu^{(d)}$ of the $\hat{p}$ modes for each dimension $d$ of the support (15). We use $\mu^{(d)}$ to compute the candidate bounds of the adapted support $\theta_{\min}^{(d)}$$^{\prime}$, $\theta_{\max}^{(d)}$$^{\prime}$ (16, LABEL: and 17). We then evaluate these candidate bounds for their physical feasibility (18, LABEL: and 22), and adjust them accordingly. The result is the adapted support.
@@ V-A Sim2Sim parameter inference setup
In all experiments, we use an MDNN for the CDF approximation; its hyperparameters are given in App. B, Table II. We perform $10$ inference iterations for Lotka-Volterra and $15$ for M/G/1 (alg. 1, 8). Each iteration augments the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled from the latest posterior.
@@ VI-B Simulation setup
Consistent with our support adaptation methodology (section IV), in all inference experiments, we consider a support $\Theta_{0}$, from which we initially sample the system parameters, and a broader domain $\Phi$ which is the physical feasibility limit of the support adaptation. Naturally, $\Theta_{0}\subseteq\Phi$, and similarly each inference iteration’s (index $i$) support is $\Theta_{i}\subseteq\Phi$. When training a policy with uniform DR ($\text{PPO-}\mathit{U}$ in section VII-C), we sample over $\Theta_{0}$. When training a policy using an inferred MoG for DR, we sample the support in which we converged by the last Real2Sim inference iteration (alg. 5, 18).
@@ VI-F Real2Sim parameter inference setup
We use the standard BayesSim algorithm [19] (alg. 1), and our EDGE variant (alg. 2), whose hyperparameters are given in App. D, Table V. In both cases, we use an MDNN for the CDF approximation. Our MDNN hyperparameters are given in App. B, Table II. As in alg. 1, 8, we approximate the posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each augmenting the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
@@ VI-I Domain randomisation setup
We implement DR by training our RL policies in vectorised environments [59] and launching parallel instances of the simulation, each with a different $\boldsymbol{\theta}$, sampled by the current domain distribution $p(\boldsymbol{\theta})$ [46]. $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (alg. 5, 18). For more Sim2Real robustness and to minimise the need for camera calibration, we introduce a small randomness in the simulated camera and stack placement position. For the sim camera, we uniformly sample $(x,y,z)$ offsets in $(\pm 0.025,\,\pm 0.025,\,\pm 0.025)\,$\mathrm{m}$\$, and for the stack position, we uniformly sample $\pm 0.02\,$\mathrm{m}$\$, $\pm 5\,$\mathrm{mm}$\$ $x$ and $y$ axis offsets.
@@ VII Real2Sim2Real DLO whipping performance
Figure 11: Progression of prior samples (left) and their respective MoG posteriors (right, see alg. 1, 10 and alg. 2, 10) along $15$ inference iterations for DLO-0, DLO-1, DLO-2 and DLO-3. On prior samples scatterplots, we see the progression of each inference iteration’s samples. Colourmapping indicates the iteration each sample was drawn, the lighter the colour, the later the iteration. Each iteration samples’ bounding box is plotted in dashed lines of the same colour. The bigger circular points indicate the accumulated dataset’s mean per inference iteration. A black dashed line marks its trajectory. We plot and annotate every second accumulated mean. On MoG heatmaps, component means are displayed in blue crosses and colourbars quantify likelihood. MoG progression is annotated with arrows pointing to the position of each iteration’s weighted component mean. We annotate every second weighted component mean. For coherence the heatmaps show only the result of the last iterations (also in fig. 12).

### v2

@@ II-B Iterative posterior refinement
In addition to iterative posterior refinement, sequential inference can involve progressive densification of the dataset $\mathcal{D}$ used to train the CDF $q_{\phi}$ (alg. 1, line 10[35]. In this way, our belief about $\boldsymbol{\theta}$ evolves recursively as more informative simulated evidence accumulates. When the samples drawn from each iteration’s posterior are aggregated in the same dataset, densifying certain areas of the supported parameter space, this proxies the reweighing of the prior to adaptively shift focus toward informative regions. Thus, in practice, we can ignore the proposal prior $\tilde{p}(\boldsymbol{\theta})$ and infer the posterior using only the (desirable) prior $p(\boldsymbol{\theta})$ and the CDF.
@@ III A demo of the support misspecification issue
We experiment with the stochastic dynamical benchmarks of the Lotka-Volterra and M/G/1 queue, as implemented for the seminal LFI work of [11]. The hyperparameters of the MDNN used to approximate the CDF (alg. 1, line 12) in the following experiments are given in App. B, Table II.
@@ IV Heuristic support optimisation methods
We treat support expansion as window stretching [40] rather than window sliding [9]: instead of shifting the sampling window, we iteratively accumulate new parameter samples guided by the latest posterior (alg. 1, line 10). We compactly denote a distribution $p$ with support $\Theta\subseteq\mathbb{R}^{(D)}$ as $p_{(\Theta)}$, but do this sparingly. This arbitrarily denotes that the distribution is supported by a $D$-dim. rectangular range defined by the bounds $\Theta=[\boldsymbol{\theta}_{\min}^{(d)},\boldsymbol{\theta}_{\max}^{(d)}]_{d=1}^{D}$.
@@ IV-A BayesSim-EDGE: expanding bounds via edge mass
EDGE is configurable through its hyperparameters. These are the edge mass threshold $\tau$, the edge zone fraction $\delta$, the edge expansion factor $\eta$, and the physically feasible domain limits $\Phi$. Specifically, $\delta$ is the size of the area within which we consider any accumulated probability mass to be near the edge. This is the sensitivity of our expansion criterion, which compares the accumulated probability mass with a threshold $\tau$. We denote the left and right edge zones ($\min$ and $\max$ side) as $\Delta_{L}^{(d)}$ and $\Delta_{R}^{(d)}$ for a dimension $d$ of $\boldsymbol{\theta}$, and the respective accumulated masses as $M_{L}^{(d)}$ and $M_{R}^{(d)}$. If $\tau$ has been exceeded and if an expansion would still be within $\Phi$ (lines 20, LABEL: and 23), we consider the expansion criterion to be satisfied and expand the respective domain boundary by a factor of $\eta$ (lines 21, LABEL: and 24). If $\tau$ has been exceeded, but the expansion would have been beyond the extremum of the respective $\Phi$ limit, then this becomes the new boundary.
@@
EDGE explores beyond the initial parameter ranges if the inferred posterior (line 12) confidently points toward the edges of the domain. This confidence is tuned through $\tau$ and $\delta$. Thus, within $N_{\text{LFI}}$ iterations, we can converge to an adapted support $\Theta_{N_{\text{LFI}}}$ in conjunction with a posterior $\hat{p}$.
@@ IV-B BayesSim-MODE: mode-shift-based expansion
Similarly to EDGE, the standard BayesSim assumptions apply. We also have MODE’s hyperparameters, which are the mode shift threshold $\nu_{\text{TH}}$, the bounds proximity threshold $\rho$, the MoG weights summary threshold $\tau$, the bounds expansion factor $\eta$, and the physically feasible domain limits $\Phi$. The main difference between MODE and EDGE is that we now calculate the shift of the MoG posterior modes $\boldsymbol{\mu}_{i}$ compared to the previous iteration $\boldsymbol{\mu}_{i-1}$ (line 21) along with their normalised proximity to the bounds (line 23). Thus, we do not check for $\Theta$ adaptation at the end of the first inference iteration. In all subsequent iterations, the MoG details of the previous iteration are accessible through the reference prior.
We accumulate the mixture coefficients (weights) of each mode that satisfies the shift and proximity thresholds for the left or right bound (lines 26, LABEL: and 28) of a parameter $d$. The accumulated weights $W_{L}^{(d)}$, $W_{R}^{(d)}$ are used for the expansion criteria (lines 31, LABEL: and 34). If $\tau$ has been exceeded and if an expansion would be within $\Phi$ (lines 31, LABEL: and 34), we expand it by a factor of $\eta$ (lines 32, LABEL: and 35). If $\tau$ has been exceeded, but the expansion would have been beyond the corresponding $\Phi$ limit, then this becomes the new boundary.
@@ IV-C BayesSim-CENTRE: centring-based adaptive bounds
Thus, beyond the standard BayesSim assumptions, we now only have physically feasible domain limits $\Phi$. In each inference iteration, following the posterior $\hat{p}$ update (line 10), we compute the weighed mean $\mu^{(d)}$ of the $\hat{p}$ modes for each dimension $d$ of the support (line 15). We use $\mu^{(d)}$ to compute the candidate bounds of the adapted support $\theta_{\min}^{(d)}$$^{\prime}$, $\theta_{\max}^{(d)}$$^{\prime}$ (lines 16, LABEL: and 17). We then evaluate these candidate bounds for their physical feasibility (lines 18, LABEL: and 22), and adjust them accordingly. The result is the adapted support.
@@ V-A Sim2Sim parameter inference setup
In all experiments, we use an MDNN for the CDF approximation; its hyperparameters are given in App. B, Table II. We perform $10$ inference iterations for Lotka-Volterra and $15$ for M/G/1 (alg. 1, line 8). Each iteration augments the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled from the latest posterior.
@@ VI-B Simulation setup
Consistent with our support adaptation methodology (section IV), in all inference experiments, we consider a support $\Theta_{0}$, from which we initially sample the system parameters, and a broader domain $\Phi$ which is the physical feasibility limit of the support adaptation. Naturally, $\Theta_{0}\subseteq\Phi$, and similarly each inference iteration’s (index $i$) support is $\Theta_{i}\subseteq\Phi$. When training a policy with uniform DR ($\text{PPO-}\mathit{U}$ in section VII-C), we sample over $\Theta_{0}$. When training a policy using an inferred MoG for DR, we sample the support in which we converged by the last Real2Sim inference iteration (alg. 5, line 18).
@@ VI-F Real2Sim parameter inference setup
We use the standard BayesSim algorithm [19] (alg. 1), and our EDGE variant (alg. 2), whose hyperparameters are given in App. D, Table V. In both cases, we use an MDNN for the CDF approximation. Our MDNN hyperparameters are given in App. B, Table II. As in alg. 1, line 8, we approximate the posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each augmenting the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
@@ VI-I Domain randomisation setup
We implement DR by training our RL policies in vectorised environments [59] and launching parallel instances of the simulation, each with a different $\boldsymbol{\theta}$, sampled by the current domain distribution $p(\boldsymbol{\theta})$ [46]. $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (alg. 5, line 18). For more Sim2Real robustness and to minimise the need for camera calibration, we introduce a small randomness in the simulated camera and stack placement position. For the sim camera, we uniformly sample $(x,y,z)$ offsets in $(\pm 0.025,\,\pm 0.025,\,\pm 0.025)\,$\mathrm{m}$\$, and for the stack position, we uniformly sample $\pm 0.02\,$\mathrm{m}$\$, $\pm 5\,$\mathrm{mm}$\$ $x$ and $y$ axis offsets.
@@ VII Real2Sim2Real DLO whipping performance
Figure 11: Progression of prior samples (left) and their respective MoG posteriors (right, see alg. 1, line 10 and alg. 2, line 10) along $15$ inference iterations for DLO-0, DLO-1, DLO-2 and DLO-3. On prior samples scatterplots, we see the progression of each inference iteration’s samples. Colourmapping indicates the iteration each sample was drawn, the lighter the colour, the later the iteration. Each iteration samples’ bounding box is plotted in dashed lines of the same colour. The bigger circular points indicate the accumulated dataset’s mean per inference iteration. A black dashed line marks its trajectory. We plot and annotate every second accumulated mean. On MoG heatmaps, component means are displayed in blue crosses and colourbars quantify likelihood. MoG progression is annotated with arrows pointing to the position of each iteration’s weighted component mean. We annotate every second weighted component mean. For coherence the heatmaps show only the result of the last iterations (also in fig. 12).